Grade 12th passPhysical Chemistry

water is electrolysed and 33.6ml h2+o2 gas is obtained at stp ten find out charged flow through circuit

naimatullah khan
8 Years agoGrade 12th pass

1 Answer

Rituraj Tiwari
5 Years ago

To find the charge that has flowed through the circuit during the electrolysis of water, we need to first understand the process and the relationship between the volume of gases produced and the amount of electric charge involved. The electrolysis of water involves splitting water (H2O) into hydrogen (H2) and oxygen (O2) gases using an electric current.

Understanding Electrolysis and Gas Production

During electrolysis, water decomposes into hydrogen and oxygen gases. The balanced chemical equation for the electrolysis of water is:

  • 2 H2O (l) → 2 H2 (g) + O2 (g)

This indicates that for every 2 moles of water that are electrolyzed, 2 moles of hydrogen and 1 mole of oxygen are produced.

Volume of Gases at STP

At Standard Temperature and Pressure (STP), 1 mole of any gas occupies 22.4 liters (or 22,400 milliliters). Given that you have obtained 33.6 mL of gas, we need to determine how many moles of gas this corresponds to:

  • 33.6 mL = 0.0336 L
  • Number of moles of gas = Volume (L) / Molar volume (L/mol) = 0.0336 L / 22.4 L/mol = 0.0015 moles

Determining Moles of Each Gas

Since the reaction produces 2 moles of hydrogen for every 2 moles of water and 1 mole of oxygen, we can distribute the moles of gas produced:

  • For hydrogen (H2): 2/3 of the total moles = (2/3) * 0.0015 = 0.0010 moles
  • For oxygen (O2): 1/3 of the total moles = (1/3) * 0.0015 = 0.0005 moles

Calculating Charge Flowed Through the Circuit

The charge (Q) can be calculated using Faraday's law of electrolysis, which states that the amount of substance produced at an electrode during electrolysis is directly proportional to the total electric charge passed through the electrolyte. The formula for charge is:

  • Q = n × F

Where:

  • Q is the total charge in coulombs (C)
  • n is the number of moles of electrons transferred
  • F is Faraday's constant, approximately 96485 C/mol

For the electrolysis of water, 2 moles of electrons are needed to produce 1 mole of H2 and 4 moles of electrons are needed to produce 1 mole of O2. Thus, for our case:

  • For 0.0010 moles of H2, electrons needed = 0.0010 moles × 2 = 0.0020 moles of electrons
  • For 0.0005 moles of O2, electrons needed = 0.0005 moles × 4 = 0.0020 moles of electrons

Total Moles of Electrons

In total, you need to account for the electrons involved in producing both gases:

  • Total moles of electrons = 0.0020 + 0.0020 = 0.0040 moles of electrons

Final Calculation of Charge

Now, substituting into the equation for charge:

  • Q = 0.0040 moles × 96485 C/mol = 385.94 C

Therefore, the total charge that has flowed through the circuit during the electrolysis of water to produce 33.6 mL of gas at STP is approximately 386 coulombs.